OpenStudy (anonymous):

Solve x^2 + 9x + 9 = 0.

OpenStudy (anonymous):

Possible intermediate steps: x^2+9 x+9 = 0 Subtract 9 from both sides: x^2+9 x = -9 Add 81/4 to both sides: x^2+9 x+81/4 = 45/4 Factor the left hand side: (x+9/2)^2 = 45/4 Take the square root of both sides: sqrt(x+9/2) = (3 sqrt(5))/2 Eliminate the absolute value: x+9/2 = -(3 sqrt(5))/2 or x+9/2 = (3 sqrt(5))/2 Subtract 9/2 from both sides: x = -3/2 (3+sqrt(5)) or x+9/2 = (3 sqrt(5))/2 Subtract 9/2 from both sides: x = -3/2 (3+sqrt(5)) or x = 3/2 (-3+sqrt(5))

OpenStudy (alfie):

There you go.

OpenStudy (anonymous):

Thank you.